OpenStudy (anonymous):

Simplify 6(2r − 5) + 2r a 4r − 30 b 4r + 30 c 14 r − 30 d 12 r − 30

OpenStudy (ny,ny):

6(2r-5) +2r 12r-30+2r 14r-30

OpenStudy (anonymous):

okay thank you... can you help me with this one? Your Math class has 32 students. You would like to treat the class to cookies. You would like each person to have 5 cookies. Which expression would be a quick way to figure out how many cookies you would need to bake? a 5(32 + 2) b 5(30 + 2) c 5(32 − 2) d 5(30 − 2)

OpenStudy (anonymous):

is it b?

OpenStudy (ny,ny):

hmm. to me the easiest way to do this is 32 x 5. which is 160. i also thought it was b but when i simplified i get 152.

OpenStudy (anonymous):

Can I help here?

OpenStudy (ny,ny):

yeah go ahead.

OpenStudy (anonymous):

what are you guys stuck on?

OpenStudy (mathstudent55):

b is correct. 5(30 + 2) = 5(32) = 160

OpenStudy (ny,ny):

OMG IM SO STUPID.

OpenStudy (anonymous):

to check to see take the orgainal problem and times the 32 to 5 and see if you get 160

OpenStudy (anonymous):

so yes b is correct

OpenStudy (mathstudent55):

You can also do it using the distributive property: 5(30 + 2) = 5(30) + 5(2) = 150 + 10 = 160

OpenStudy (anonymous):

that is the other way too.. :)

OpenStudy (anonymous):

thanks guys appreciate it

OpenStudy (anonymous):

not a problem and have a good night or day wherever you are..

OpenStudy (ny,ny):

night :P

OpenStudy (anonymous):

Hey NY NY I am now your fan :)

OpenStudy (ny,ny):

me too :)